### v2

# Density estimation via mixture discrepancy and moments

Abstract
With the aim of generalizing histogram statistics to higher dimensional cases, density estimation via discrepancy based sequential partition (DSP) has been proposed to learn an adaptive piecewise constant approximation defined on a binary sequential partition of the underlying domain, where the star discrepancy is adopted to measure the uniformity of particle distribution. However, the calculation of the star discrepancy is NP-hard and it does not satisfy the reflection invariance and rotation invariance either. To this end, we use the mixture discrepancy and the comparison of moments as a replacement of the star discrepancy, leading to the density estimation via mixture discrepancy based sequential partition (DSP-mix) and density estimation via moment-based sequential partition (MSP), respectively. Both DSP-mix and MSP are computationally tractable and exhibit the reflection and rotation invariance. Numerical experiments in reconstructing Beta mixtures, Gaussian mixtures and heavy-tailed Cauchy mixtures up to 30 dimension are conducted, demonstrating that MSP can maintain the same accuracy compared with DSP, while gaining an increase in speed by a factor of two to twenty for large sample size, and DSP-mix can achieve satisfactory accuracy and boost the efficiency in low-dimensional tests ($d\leq 6$), but might lose accuracy in high-dimensional problems due to a reduction in partition level.
keywords:
Density estimation; Adaptive sequential partition; Discrepancy; Hausdorff moment problem; Kullback-Leibler divergence; Hellinger distance; Beta mixtures; Gaussian mixtures; Cauchy mixtures
1 Introduction
As a fundamental problem in statistics, density estimation aims at constructing an estimate of their common density function $q(\bm{y})$ for given $N$ observations $\bm{y}_{1},\bm{y}_{2},\dots,\bm{y}_{N}$. It has various applications in uncertainty quantification [1], Bayesian inference [2] and scientific computing [3, 4, 5, 6]. A simple and classical approach is the histogram statistics as a uniform piecewise constant density estimator. Unfortunately, it is rather difficult to apply the histogram to high-dimensional problems as both bin size and required sample size grow exponentially in dimension, which is well known as the curse of dimensionality (CoD) [7, 3].
This paper focuses on nonparametric approaches for multivariate density estimation, especially for the case when the dimension $d$ is moderately large, say, 6306–30 [8, 9]. The benefit of nonparametric methods is their ability to achieve estimation optimality for any input distribution as more data are observed. A widely used method is the kernel density estimation (KDE), treating each data point as the center of a kernel function like the Gaussian functions and smoothed splines [7]. But the accuracy of the kernel estimator becomes very sensitive to the choice of the window size and the shape of the kernel [10], and the magic lies in how to balance the bias and variance of the estimator, like using the smoothing properties of linear diffusion processes [11] or choosing adaptive bandwidths [12, 10]. Another approach is based on tensor decompositions, with huge advantages in that the computational cost of the construction, the storage requirements and the operations required for conditional distribution method sampling from the distributional approximation all scale linearly with dimension [4, 5, 13]. However, it still relies heavily on the low-rank assumption of the underlying density, whereas the detection and characterization of local features of multivariate density estimation are usually prohibitive [8]. Machine learning approaches have recently been drawing a growing attention, e.g., the density estimation through normalizing flows [14] and deep generative neural networks [15]. In practice, they require extensive tuning to perform well [16].
Here we mainly discuss a data-driven density estimation method, termed tree-based density estimation. This features a class of estimators which employs simple and flexible binary partitions to adapt to the underlying density function, along with the decision tree using stopping times in a data-driven way [2]. The decision tree based method has a great potential to overcome the burden imposed by the high dimensionality, especially when the dimension is only moderately large and the density function exhibits certain spatial features that can be leveraged of [16]. Essentially, it is equivalent to cluster the samples into small nonintersecting sets, each supported by a tree-structured density [17]. An adaptive nonparametric density estimation can be constructed by either the Bayesian Sequential Partition (BSP) [8, 16] or the Discrepancy based Sequential Partition (DSP) [18]. Compared with BSP, the computational cost of DSP is much cheaper due to its greedy construction, and is fully capable to handle higher dimension [6].
To be more specific, DSP uses the star discrepancy, a concept from the quasi-Monte Carlo method, to measure the uniformity of the observations [19, 20]. For a point set $\left(\bm{y}_{1},\dots,\bm{y}_{n}\right)\subset\left[0,1\right]^{d}$, its star discrepancy reads
$D^{*}\left(\bm{y}_{1},\ldots,\bm{y}_{n}\right)=\sup_{\boldsymbol{u}\in[0,1]^{d}}\left|\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}_{[0,\boldsymbol{u})}\left(\bm{y}_{i}\right)-\operatorname{vol}([0,\boldsymbol{u}))\right|,$ (1)
and characterizes the uniformity of the points by using the infinity norm of the difference between the cumulative distribution function of the discrete point measure $\frac{1}{n}\sum_{i=1}^{n}\delta_{\bm{y}_{i}}$ and that of the uniform measure. When the star discrepancy of the observations within a sub-domain is small, it can be reasonable to think that their distribution approaches the uniform distribution and thus the corresponding probability density can be regarded as approximately constant in the sub-domain.
The major bottleneck of DSP lies in the complexity of calculating the star discrepancy of pointsets in the $d$-dimensional unit cube, which has been proved to be an NP-hard problem [21]. As a consequence, the external approximate solver [22, 23, 24] is mandatory, but is rather time-consuming when the sample size $N$ is large. Besides, the star discrepancy lacks both reflection and rotation invariance [20, 25, 26] since the origin $\boldsymbol{0}$ is distinctive in the definition (1) of the star discrepancy. In view of such drawbacks, this paper contributes two new uniformity measurement strategies for density estimation, both of which seek a relaxation of DSP. One is to adopt the mixture discrepancy, another kind of discrepancy [20, 25, 26]. For a point set $\left(\bm{y}_{1},\dots,\bm{y}_{n}\right)\subset\left[0,1\right]^{d}$, its mixture discrepancy has an analytical expression,
$\left(D^{\text{mix}}(\mathbf{y}_{1},\ldots,\mathbf{y}_{n})\right)^{2}=$ $\left(\frac{19}{12}\right)^{d}-\frac{2}{n}\sum_{i=1}^{n}\prod_{j=1}^{d}\left(\frac{5}{3}-\frac{1}{4}\left|y_{ij}-\frac{1}{2}\right|-\frac{1}{4}\left|y_{ij}-\frac{1}{2}\right|^{2}\right)$ (2)
$+\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{k=1}^{n}\prod_{j=1}^{d}\left(\frac{15}{8}-\frac{1}{4}\left|y_{ij}-\frac{1}{2}\right|-\frac{1}{4}\left|y_{kj}-\frac{1}{2}\right|\right$.
$\left.-\frac{3}{4}\left|y_{ij}-y_{kj}\right|+\frac{1}{2}\left|y_{ij}-y_{kj}\right|^{2}\right),$
where $\bm{y}_{i}=(y_{i1},\dots,y_{id})$. It is thus easy to calculate, and possesses good theoretical properties. The interested readers are referred to [20, 25, 26] for more details about the mixture discrepancy including the motivation and the calculation. We use the same algorithm framework as DSP except for replacing the star discrepancy with the mixture discrepancy. The resulting method is dubbed the DSP-mix method.
The other is the density estimation via moment-based sequential partition (MSP), which still adopts the partitioning framework in DSP, but the way to measure the uniformity of the observations is replaced by comparing the gap in moments between the particle distribution and the uniform density. Whether the density function can be uniquely determined by the moments of all orders is the classical moment problem in probability theory [27, 28, 29]. For the case where the probability density is defined on a compact set, this problem is called the Hausdorff moment problem, and the answer is positive [30]. When the moments of truncated orders of the observations on the compact set are close to those of the uniform distribution, we can consider that the observations are approximately uniformly distributed. Numerical experiments in reconstructing Beta mixtures, Gaussian mixtures and heavy-tailed Cauchy mixtures up to 30 dimension are conducted, demonstrating that MSP can maintain the same accuracy compared with DSP, while gaining an increase in speed by a factor of two to twenty for large sample size, and DSP-mix can achieve satisfactory accuracy in low-dimensional tests ($d\leq 6$) and run significantly faster than DSP, but might lose accuracy in high-dimensional problems due to a reduction in partition level.
The rest of this paper is organized as follows. The DSP-mix and MSP methods are introduced in Sections 2 and 3, respectively. Section 4 conducts several numerical experiments and compares DSP-mix and MSP with DSP in performance. The paper is concluded in Section 5.
2 Discrepancy sequential partition via star discrepancy or mixed discrepancy
Without loss of generality, assume that the observations $\bm{y}_{i}\in\Omega,\,i=1,2,\dots,N$, and the computational domain $\Omega$ is decomposed into $L$ non-intersecting sub-domains,
$\Omega=\bigcup_{l=1}^{L}\Omega_{l}.$ (3)
DSP uses an adaptive piecewise constant density estimator to reconstruct the $d$-D probability density function from the observations [18],
$\hat{p}(\bm{y})=\sum_{l=1}^{L}c_{l}\mathbbm{1}_{\Omega_{l}}(\bm{y}),\quad c_{l}=\frac{1}{N|\Omega_{l}|}\sum_{i=1}^{N}\mathbbm{1}_{\Omega_{l}}(\bm{y}_{i}).$ (4)
As presented in Figure 1, DSP adopts a sequential binary partitioning strategy to obtain the decomposition (3).
Figure 1: An illustration of tree-based density estimation.
2.1 Setting of tree-based partitioning
Algorithm 1 The framework of tree-based density estimation.
Input: The domain $\Omega$, particle set $\bm{y}_{i},\,i=1,2,\dots,N$, parameters $\theta,m$.
Output: An adaptive piecewise constant approximation of probability density function $q(\bm{y})$.
[…]
2:while true do
3:  flag = 0;
4:  for all $\Omega_{l}=\prod\limits_{j=1}^{d}\left[a_{j}^{(l)},b_{j}^{(l)}\right]\in\mathcal{P}_{L}$ do
5:   Denote $S_{l}=\{(y^{(l)}_{i1},\dots,y^{(l)}_{id})\}_{i=1}^{n}$ as the subset of particles that fall within $\Omega_{l}$ and $n=|S_{l}|$;
6:   Measure the uniformity of $S_{l}$;
7:   if $S_{l}$ cannot be approximated as a uniform distribution, e.g., $D^{\ast}(\widetilde{S}_{l})>\frac{\theta\sqrt{N}}{n_{l}}$ then.
8:     $L\leftarrow L+1$;
9:     Choose a splitting node $\widetilde{s}_{i_{0},j_{0}}^{(l)}$ according to
$\widetilde{s}_{i_{0},j_{0}}^{(l)}=\underset{s_{i_{0},j_{0}}^{(l)}}{\arg\max}\left|\frac{n_{i_{0},j_{0}}}{n}-\frac{i_{0}}{m}\right|$
10:     Divide $\Omega_{l}$ into $\Omega_{l}^{(1)}\cup\Omega_{l}^{(2)}$ as
$\Omega_{l}^{(1)}=\prod_{j=1}^{j_{0}-1}\left[a_{j}^{(l)},b_{j}^{(l)}\right]\times\left[a_{j_{0}}^{(l)},s_{i_{0},j_{0}}^{(l)}\right]\times\prod_{j=j_{0}+1}^{d}\left[a_{j}^{(l)},b_{j}^{(l)}\right],\quad\Omega_{l}^{(2)}=\Omega_{l}\backslash\Omega_{l}^{(1)}$
11:     $\Omega_{l}\leftarrow\Omega_{l}^{(1)},\Omega_{L+1}\leftarrow\Omega_{l}^{(2)}$;
12:     flag = 1;
13:     Break;
14:   end if
15:  end for
16:  if flag = 1 then
17:   $\mathcal{P}_{L+1}\leftarrow\left\{\Omega_{1},\ldots,\Omega_{L+1}\right\}$;
18:  else
19:   Break;
20:  end if
21:end while
22:Calculate a piecewise constant density estimator $\hat{p}(\bm{y})$ corresponding to the partition $\mathcal{P}_{L}$ by
$\hat{p}(\bm{y})=\sum_{l=1}^{L}c_{l}\mathbbm{1}_{\Omega_{l}}(\bm{y}),\quad c_{l}=\frac{1}{N|\Omega_{l}|}\sum_{i=1}^{N}\mathbbm{1}_{\Omega_{l}}(\bm{y}_{i})$
When the star discrepancy given in Eq. (1) of the observations within the sub-domain $\Omega_{l}$ is less than a predetermined threshold, it can be approximately considered that the observations within this sub-domain are uniformly distributed, and the corresponding density function can be approximated as a constant. Algorithm 1 presents the detailed process. It can be readily seen there that if there exists a sub-domain $\Omega_{l}$ that fails to pass the uniformity test (see Lines 7-14), then $\Omega_{l}$ is split along a certain coordinate and two new sub-domains, $\Omega_{l}^{(1)}$ and $\Omega_{l}^{(2)}$, are generated. Subsequently, these two newly generated sub-domains are combined with the remaining sub-domains to update the sub-domain set $\mathcal{P}_{L+1}$ (see Line 17).
Suppose $\Omega_{l}=\prod\limits_{j=1}^{d}\left[a_{j}^{(l)},b_{j}^{(l)}\right]$ and $S_{l}=\{(y^{(l)}_{i1},\dots,y^{(l)}_{id})\}_{i=1}^{n}$ to be the subset of particles that fall within $\Omega_{l}$. Two key issues are left to be specified in Algorithm 1: Where to split and whether to split.
  • 1.
    Where to split: We adopt m1m-1 dividing points along each coordinate, and choose the uniformly distributed points for simplicity: $s_{i_{0},j_{0}}^{(l)}=a_{j_{0}}^{(l)}+\frac{i_{0}}{m}\left(b_{j_{0}}^{(l)}-a_{j_{0}}^{(l)}\right)$ with $i_{0}=1,2,\dots,m-1$ and $j_{0}=1,2,\dots,d$. Accordingly, DSP splits $\Omega_{l}$ using the dividing point $\widetilde{s}_{i_{0},j_{0}}^{(l)}$ that divides $\Omega_{l}$ into $\Omega_{l}^{(1)}$ and $\Omega_{l}^{(2)}$. Here $\widetilde{s}_{i_{0},j_{0}}^{(l)}$ is selected according to
    $\widetilde{s}_{i_{0},j_{0}}^{(l)}=\underset{s_{i_{0},j_{0}}^{(l)}}{\arg\max}\left|\frac{n_{i_{0},j_{0}}}{n}-\frac{i_{0}}{m}\right|,$ (5)
    where $n_{i_{0},j_{0}}$ count the number of particles in $S_{l}$ that fall within $\Omega_{l}^{(1)}$, and
    $\Omega_{l}^{(1)}=\prod_{j=1}^{j_{0}-1}\left[a_{j}^{(l)},b_{j}^{(l)}\right]\times\left[a_{j_{0}}^{(l)},s_{i_{0},j_{0}}^{(l)}\right]\times\prod_{j=j_{0}+1}^{d}\left[a_{j}^{(l)},b_{j}^{(l)}\right],\quad\Omega_{l}^{(2)}=\Omega_{l}\backslash\Omega_{l}^{(1)}.$ (6)
    The intuition behind Eq. (5) is to find the most non-uniform sub-domain $\Omega_{l}^{(1)}$ for a given set of dividing points $\{s_{i_{0},j_{0}}^{(l)}\big|i_{0}=1,2,\dots,m-1,j_{0}=1,2,\dots,d\}$;
  • 2.
    Whether to split: The linear scaling $\Omega_{l}\rightarrow\left[0,1\right]^{d}$ is performed first before using the discrepancy as the uniformity measurement. We scale $S_{l}$ to $\widetilde{S}_{l}=\{(\frac{y^{(l)}_{i1}-a_{1}^{(l)}}{b_{1}^{(l)}},\dots,\frac{y^{(l)}_{id}-a_{d}^{(l)}}{b_{d}^{(l)}})\}_{i=1}^{n}$. The partitioning terminates when the discrepancy of particles that fall within $\Omega_{l}$ satisfying
    $D^{\ast}(\widetilde{S}_{l})\leq\frac{\theta\sqrt{N}}{n_{l}}.$ (7)
    The parameter $\theta$ governs the depth of partition. The smaller $\theta$ is, the finer the partitioning it induces.
2.2 Theoretical analysis of DSP
We can use the $\ast$-total variation distance to measure the difference between the density estimator $\hat{p}(\bm{y})$ and the empirical measure $\mu=\frac{1}{N}\sum_{i=1}^{N}\delta_{\bm{y}_{i}}$. For two measures $\mu$ and $\nu$, the $\ast$-total variation distance is defined as
[…]
where $\mu(A)=\int_{\Omega}1_{A}(\bm{y})\textup{d}\mu(\bm{y})$. A rough error estimator is given as follows.
Theorem 2.1.
For the partition $\Omega=\cup_{l=1}^{L}\Omega_{l}$, suppose $D^{*}(\widetilde{S}_{l})\leq\theta\sqrt{N}/n_{l}$ in each subregion $\Omega_{l}$, then it has
$D^{\ast}[\mu,\hat{p}]\leq\frac{L\theta}{\sqrt{N}}.$ (8)
Thus for any function $f$ with bounded variation $V_{HK}(f;\Omega)$ in the sense of Hardy and Krause, the transport cost has an upper bound
$\Big|\int_{\Omega}f\hat{p}\textup{d}\bm{y}-\int_{\Omega}f\textup{d}\mu\Big|\leq\frac{L\theta}{\sqrt{N}}V_{HK}(f;\Omega).$ (9)
Proof.
Denote by $\{\bm{y}_{i}\in A\}$ the count of particles in $A$, and $\bm{y}_{i}^{(l)}$ the particles in the subregion $\Omega_{l}$. For arbitrary $A\subset\Omega$, it has
$\mu(A)=\frac{1}{N}\#\{\bm{y}_{i}\in A\},\quad\hat{p}(A)=\sum\limits_{l=1}^{L}\frac{n_{l}}{N}\frac{\textup{vol}(A\cap\Omega_{l})}{\textup{vol}(\Omega_{l})}.$
Since $A\cap\Omega_{l}$ is still anchored at the left corner of $\Omega_{l}$, it yields that
$\begin{split}|\mu(A)-\nu(A)|&=\left|\frac{1}{N}\#\{\bm{y}_{i}\in A\}-\sum_{l=1}^{L}\frac{n_{l}}{N}\frac{\textup{vol}(A\cap\Omega_{l})}{\textup{vol}(\Omega_{l})}\right|\\ &=\left|\frac{1}{N}\sum_{l=1}^{L}\#\{\bm{y}_{i}^{(l)}\in A\cap\Omega_{l}\}-\sum_{l=1}^{L}\frac{n_{l}}{N}\frac{\textup{vol}(A\cap\Omega_{l})}{\textup{vol}(\Omega_{l})}\right|\\ &\leq\frac{1}{N}\sum_{l=1}^{L}n_{l}\left|\frac{\#\{\bm{y}_{i}^{(l)}\in A\cap\Omega_{l}\}}{n_{l}}-\frac{\textup{vol}(A\cap\Omega_{l})}{\textup{vol}(\Omega_{l})}\right|\\ &\leq\frac{1}{N}\sum_{l=1}^{L}n_{l}D^{*}(\widetilde{S}_{l})\leq\frac{1}{N}\sum_{l=1}^{L}n_{l}\frac{\theta\sqrt{N}}{n_{l}}=\frac{L\theta}{\sqrt{N}}.\end{split}$
By taking supremum over $A\subset\Omega$, it arrives at $D^{*}[\mu,\hat{p}]\leq\frac{L\theta}{\sqrt{N}}$. Eq. (9) is a direct consequence of the Koksma-Hlawka inequality. ∎
Since the partition level $L$ actually depends on $\theta$ and $N$, the above rough bound is useful only when $L$ is not so large. In particular, the total variation $V_{HK}(f;\Omega)$ scales linearly with $d$. For $f(\bm{y})=\sum_{i=1}^{d}y_{i}^{2}$ or $f(\bm{y})=\sum_{j\neq i}\sum_{i=1}^{d}y_{i}y_{j}$, $V_{HK}(f;\Omega)$ scales quadratically with $d$. As a result, Eq. (9) controls the discrepancy in the first and second moments of $\hat{p}$ and $\mu$.
2.3 Relaxation by the mixed discrepancy
Despite that there are some efficient heuristic algorithms for calculating the star discrepancy, like the TA-improved method [23] and the ODE annealing approach [24], the computational cost dramatically grows as sample size $N$ and dimension $d$ increase. Apart from the huge computational complexity, the star discrepancy possesses two drawbacks as follows [25].
(a) $D^{*}=\frac{23}{48}$, $D^{\text{mix}}=\frac{2719}{18432}$.
(b) $D^{*}=\frac{9}{16}$, $D^{\text{mix}}=\frac{2719}{18432}$.
Figure 2: Reflection invariance test, $d=2,N=3$: Figures 2(a) and 2(b) are symmetric about $x_{1}=0.5$. Intuitively, the reflection transformation should not change the uniformity of points. However, the star discrepancies of Figures 2(a) and 2(b) are different.
(a) $D^{*}=\frac{13}{24},D^{\text{mix}}=\frac{301}{1024}$.
(b) $D^{*}=\frac{7}{16},D^{\text{mix}}=\frac{301}{1024}$.
Figure 3: Rotation invariance test, $d=2,N=3$: Figure 3(b) is obtained by rotating Figure 3(a) 90 degrees clockwise. Intuitively, rotation should not change the uniformity of the points. However, the star discrepancies of Figure 3(a) and 3(b) are different.
  1. (1)
    The star discrepancy does not satisfy reflection invariance [25, 26]: The point sets in Figures 2(a) and 2(b) are symmetric about the line $x_{1}=0.5$, but the values of their star discrepancies are not equal, which is counterintuitive. The reason for this phenomenon lies in the fact that the origin $\boldsymbol{0}$ is distinctive in the definition of the star discrepancy;
  2. (2)
    The star discrepancy does not satisfy rotation invariance [25, 26]: Figure 3(b) is obtained by rotating Figure 3(a) 90 degrees clockwise, and the value of the star discrepancy changes after the rotation. The reason for this phenomenon is also that the origin $\boldsymbol{0}$ is distinctive in the definition of the star discrepancy.
To fulfill reflection and rotation invariances, we adopt the mixed discrepancy (2) as introduced in [20, 25, 26]. The resulting DSP-mix inherits the algorithm framework of DSP and only differs in the measure of uniformity, where the mixture discrepancy $D^{\text{mix}}(\widetilde{S}_{l})$ to adopted to measure the uniformity of $S_{l}$ in Line 6 of Algorithm 1,
$D^{\text{mix}}(\widetilde{S}_{l})\leq\frac{\theta\sqrt{N}}{n_{l}}.$ (10)
Compared with the star discrepancy adopted in DSP, the mixture discrepancy has the following advantages [26].
  1. (1)
    Easy to calculate: The mixture discrepancy can be accurately calculated with a computational complexity of $\mathcal{O}(n^{2}d)$ and thus any external solver is not required. Hence DSP-mix is expected to handle the density estimation problems more efficiently than DSP. In [31], DSP-mix is used to efficiently reconstruct the values of the Coulomb collision term from particles;
  2. (2)
    Reflection invariance: The mixture discrepancy satisfies the reflection invariance [26]. Suppose that $\left(\widetilde{\bm{y}}_{1},\dots,\widetilde{\bm{y}}_{n}\right)$ is the points obtained by reflecting $\left(\bm{y}_{1},\dots,\bm{y}_{n}\right)$ about the plane $y_{j_{0}}=1/2$, then
    $\widetilde{\bm{y}}_{i}=\left(y_{i1},\dots,y_{i,j_{0}-1},1-y_{i,j_{0}},y_{i,j_{0}+1},\dots,y_{id}\right),\quad i=1\dots,n,$ (11)
    and $\left(D^{\text{mix}}(\widetilde{\bm{y}}_{1},\dots,\widetilde{\bm{y}}_{n}\right))^{2}=\left(D^{\text{mix}}(\bm{y}_{1},\ldots,\bm{y}_{n}\right))^{2}$. For the two point sets in Figures 2(a) and 2(b), their mixture discrepancies are both $\frac{2719}{18432}$;
  3. (3)
    Rotation invariance: The mixture discrepancy satisfies the rotation invariance [26]. Suppose that $\left(\widetilde{\bm{y}}_{1},\dots,\widetilde{\bm{y}}_{n}\right)$ is the points obtained by rotating $\left(\bm{y}_{1},\dots,\bm{y}_{n}\right)$ 90 degrees clockwise about the $y_{i_{0}}-y_{j_{0}}$ plane, then for $\forall i=1,\dots,n$,
    $\widetilde{y}_{ij}=\begin{cases}y_{ij},&j\neq i_{0},j_{0},\\ \cos\left(-\pi/2\right)(y_{i,i_{0}}-0.5)-\sin\left(-\pi/2\right)(y_{i,j_{0}}-0.5)+0.5=y_{i,j_{0}},&j=i_{0},\\ \sin\left(-\pi/2\right)(y_{i,i_{0}}-0.5)+\cos\left(-\pi/2\right)(y_{i,j_{0}}-0.5)+0.5=1-y_{i,i_{0}},&j=j_{0},\end{cases}$
    and it can be easily proved that $(D^{\text{mix}}\left(\widetilde{\bm{y}}_{1},\dots,\widetilde{\bm{y}}_{n}\right))^{2}=(D^{\text{mix}}\left(\bm{y}_{1},\ldots,\bm{y}_{n}\right))^{2}$. For the two point sets in Figures 3(a) and 3(b), their mixture discrepancies are both $\frac{301}{1024}$.
3 MSP: Moment-based Sequential Partition
MSP still adopts the framework of Algorithm 1. Instead of using the discrepancy as the uniformity measurement, MSP judges whether the particle set $S_{l}$ in $\Omega_{l}$ is close to a uniform distribution by comparing whether the moments of the discrete point measure $\frac{1}{N}\sum_{i=1}^{N}\delta_{\bm{y}_{i}}$ and the uniform measure are close enough.
The motivation comes from the following fact: Suppose that the $d$-dimensional random variable $\bm{y}=(y_{1},y_{2},\dots,y_{d})$ follows the probability density $q(\bm{y})$, the $d$-dimensional index $\boldsymbol{k}=(k_{1},k_{2},\dots,k_{d})$, $\bm{y}^{\boldsymbol{k}}=y_{1}^{k_{1}}y_{2}^{k_{2}}\dots y_{d}^{k_{d}}$, $\mathbb{N}_{0}=\{0,1,2,\dots\}$, then the moment of order $|\boldsymbol{k}|=\sum\limits_{i=1}^{d}k_{i}$ of the probability density $q(\bm{y})$ is defined as
$m_{\boldsymbol{k}}(q)=\int_{\mathbb{R}^{d}}\bm{y}^{\boldsymbol{k}}q(\bm{y})\,\textup{d}\bm{y},\quad\boldsymbol{k}\in\mathbb{N}_{0}^{d}.$ (12)
We call $\{m_{\boldsymbol{k}}(q),\boldsymbol{k}\in\mathbb{N}_{0}^{d}\}$ the moment sequence. Whether the probability density on a compact set can be uniquely determined by the moment sequence is the Hausdorff moment problem in probability theory, and its answer is positive [30].
Lemma 3.1.
([30]) The probability density function defined on a compact set is uniquely determined by its moment sequence.
As a consequence, for two probability densities defined on a compact set, we can judge whether they are close by comparing their moment sequences. Detailedly, for the test function $f(\bm{y})$,
$\begin{split}&\int_{\Omega}f(\mathbf{y})\hat{p}(\mathbf{y})\mathrm{d}\mathbf{y}-\int_{\Omega}f(\mathbf{y})q(\mathbf{y})\mathrm{d}\mathbf{y}\\ =&\int_{\Omega}f(\mathbf{y})\hat{p}(\mathbf{y})\mathrm{d}\mathbf{y}-\frac{1}{N}\sum_{i=1}^{N}f(\mathbf{y}_{i})+\frac{1}{N}\sum_{i=1}^{N}f(\mathbf{y}_{i})-\int_{\Omega}f(\mathbf{y})q(\mathbf{y})\mathrm{d}\mathbf{y}\\ \leq&\sum_{l=1}^{L}\frac{n_{l}}{N}\Big|\frac{1}{\textup{vol}(\Omega_{l})}\int_{\Omega_{l}}f(\mathbf{y})\mathrm{d}\mathbf{y}-\frac{1}{n_{l}}\sum_{\mathbf{y}_{i}^{(l)}\in S_{l}}f(\mathbf{y}_{i}^{(l)})\Big|\\ &+\Big|\frac{1}{N}\sum_{i=1}^{N}f(\mathbf{y}_{i})-\int_{\Omega}f(\mathbf{y})q(\mathbf{y})\mathrm{d}\mathbf{y}\Big|.\end{split}$ (13)
Thus we can choose $f(\bm{y})=\bm{y}^{k}$ to control the moment sequence in the density estimator, while the second term depends only on the Monte Carlo sampling error.
In actual numerical implementation, we discover that conducting comparisons up to the order of $|\boldsymbol{k}|\leq 2$ is capable of attaining satisfactory numerical effects, that is, only the expectation and the covariance matrix are considered. Denote the expectation of the uniform measure on the sub-domain $\Omega_{l}$ as $\boldsymbol{\mu}^{(l)}=(\mu^{(l)}_{1},\dots,\mu^{(l)}_{d})\in\mathbb{R}^{d}$, and the covariance matrix as $\boldsymbol{\Sigma}^{(l)}=(\Sigma_{ij}^{(l)})\in\mathbb{R}^{d\times d}$. The expectation of the particle subset $S_{l}$ is $\hat{\boldsymbol{\mu}}^{(l)}=(\hat{\mu}^{(l)}_{1},\dots,\hat{\mu}^{(l)}_{d})\in\mathbb{R}^{d}$, and the covariance matrix is $\boldsymbol{\hat{\Sigma}}^{(l)}=(\hat{\Sigma}_{ij}^{(l)})\in\mathbb{R}^{d\times d}$. They have the explicit calculation expressions,
$\mu^{(l)}_{j}$ $=\frac{a_{j}^{(l)}+b_{j}^{(l)}}{2},\quad j=1,\dots,d,$ (14)
$\Sigma_{jj}^{(l)}$ $=\frac{\left(b_{j}^{(l)}-a_{j}^{(l)}\right)^{2}}{12},\quad j=1,\dots,d,\quad\textup{and}\quad\Sigma_{ij}^{(l)}$ $=0,\quad i\neq j,$
and
$\hat{\boldsymbol{\mu}}^{(l)}=\frac{1}{n}\sum_{i=1}^{n}\bm{y}_{i}^{(l)},\quad\boldsymbol{\hat{\Sigma}}^{(l)}=\frac{1}{n}\sum_{i=1}^{n}\left(\bm{y}_{i}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right)\left(\bm{y}_{i}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right)^{T}.$ (15)
MSP runs the uniformity test as follows
$\left|\mu_{j}^{(l)}-\hat{\mu}_{j}^{(l)}\right|<\theta\left(b_{j}^{(l)}-a_{j}^{(l)}\right),\quad j=1,\dots,d,$ (16)
$\left|\Sigma_{jj}^{(l)}-\hat{\Sigma}_{jj}^{(l)}\right|<\theta\left|\Sigma_{jj}^{(l)}\right|,\quad j=1,\dots,d,\quad\textup{and}\quad\left|\hat{\Sigma}_{ij}^{(l)}\right|<\theta,\quad i\neq j,$
where we choose the same tolerance parameter $\theta$ in Eq. (7) used by both DSP and DSP-mix for the sake of comparison. If Eq. (16) is satisfied, MSP terminates the sequential partition. Otherwise we consider that $S_{l}$ cannot be approximated as the uniform distribution in Line 7 of Algorithm 1. Obviously, MSP has the following advantages over DSP.
  1. (1)
    Easy to calculate: The computational complexity of the expectation is $\mathcal{O}(nd)$, and the computational complexity of the covariance matrix is $\mathcal{O}(nd^{2})$. Both of them are linear functions of the number of particles $n$;
  2. (2)
    Reflection invariance: Suppose that $\left(\widetilde{\bm{y}}_{1},\dots,\widetilde{\bm{y}}_{n}\right)$ is the points obtained by reflecting $\left(\bm{y}_{1},\dots,\bm{y}_{n}\right)$ about the plane $y_{j_{0}}=\frac{a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}}{2}$, then
    $\widetilde{\bm{y}}_{i}=\left(y_{i1},\dots,y_{i,j_{0}-1},a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{i,j_{0}},y_{i,j_{0}+1},\dots,y_{id}\right),\quad i=1\dots,n.$ (17)
    Let $\widetilde{\hat{\boldsymbol{\mu}}}^{(l)}=\frac{1}{n}\sum_{i=1}^{n}\widetilde{\bm{y}}_{i}^{(l)}$ and $\widetilde{\boldsymbol{\hat{\Sigma}}}^{(l)}=\frac{1}{n}\sum_{i=1}^{n}\left(\widetilde{\bm{y}}_{i}^{(l)}-\widetilde{\hat{\boldsymbol{\mu}}}^{(l)}\right)\left(\widetilde{\bm{y}}_{i}^{(l)}-\widetilde{\hat{\boldsymbol{\mu}}}^{(l)}\right)^{T}$, then
    $\left|\mu_{j}^{(l)}-\widetilde{\hat{\mu}}_{j}^{(l)}\right|=\begin{cases}\left|\mu_{j}^{(l)}-\hat{\mu}_{j}^{(l)}\right|,&j\neq j_{0},\\ \left|\frac{a_{j}^{(l)}+b_{j}^{(l)}}{2}-(a_{j}^{(l)}+b_{j}^{(l)})+\hat{\mu}_{j}^{(l)}\right|=\left|\mu_{j}^{(l)}-\hat{\mu}_{j}^{(l)}\right|,&j=j_{0}.\end{cases}$ (18)
    For $j\neq j_{0}$, it is obvious that $\left|\Sigma_{jj}^{(l)}-\widetilde{\hat{\Sigma}}_{jj}^{(l)}\right|=\left|\Sigma_{jj}^{(l)}-\hat{\Sigma}_{jj}^{(l)}\right|$. For $j=j_{0}$, we have
    $\left|\Sigma_{j_{0}j_{0}}^{(l)}-\widetilde{\hat{\Sigma}}_{j_{0}j_{0}}^{(l)}\right|=$ $\left|\Sigma_{j_{0}j_{0}}^{(l)}-\frac{1}{n}\sum_{i=1}^{n}\left(a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{i,j_{0}}-\frac{1}{n}\sum_{k=1}^{n}(a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{k,j_{0}})\right)^{2}\right|$ (19)
    $=$ $\left|\Sigma_{j_{0}j_{0}}^{(l)}-\frac{1}{n}\sum_{k=1}^{n}\left(y_{i,j_{0}}-\frac{1}{n}\sum_{k=1}^{n}y_{k,j_{0}}\right)^{2}\right|=\left|\Sigma_{j_{0}j_{0}}^{(l)}-\hat{\Sigma}_{j_{0}j_{0}}^{(l)}\right|.$
    For $i\neq j_{0}$, $j\neq j_{0}$ and $i\neq j$, it is patently clear that $\left|\widetilde{\hat{\Sigma}}_{ij}^{(l)}\right|=\left|\hat{\Sigma}_{ij}^{(l)}\right|$. When $i\neq j_{0}$ and $j=j_{0}$,
    $\left|\widetilde{\hat{\Sigma}}_{ij_{0}}^{(l)}\right|=$ $\left|\frac{1}{n}\sum_{k=1}^{n}\left(y_{k,i}-\frac{1}{n}\sum_{h=1}^{n}y_{h,i}\right)\left(a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{k,j_{0}}-\frac{1}{n}\sum_{h=1}^{n}(a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}-y_{h,j_{0}})\right)\right|$ (20)
    $=$ $\left|\frac{1}{n}\sum_{k=1}^{n}-\left(y_{k,i}-\frac{1}{n}\sum_{h=1}^{n}y_{h,i}\right)\left(y_{k,j_{0}}-\frac{1}{n}\sum_{h=1}^{n}y_{h,j_{0}}\right)\right|=\left|\hat{\Sigma}_{ij}^{(l)}\right|.$
    The situation with $i=j_{0}$ and $j\neq j_{0}$ is analogous. Combining Eqs. (18), (19) and (20) together, the reflection transformation has no impact on the termination criterion (16). For Figures 2(a) and 2(b), the differences between the expectations and covariance matrices of the two groups of particles and those of the uniform measure are
    $\left|\boldsymbol{\mu}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right|$ $=\left(\frac{1}{12},\frac{1}{12}\right),\quad\left|\boldsymbol{\Sigma}^{(l)}-\boldsymbol{\hat{\Sigma}}^{(l)}\right|$ $=\begin{pmatrix}\frac{1}{48}&\frac{1}{48}\\ \frac{1}{48}&\frac{1}{48}\end{pmatrix}$, (21)
    where the absolute is performed element-wise;
  3. (3)
    Rotation invariance: Suppose that $\left(\widetilde{\bm{y}}_{1},\dots,\widetilde{\bm{y}}_{n}\right)$ is the points obtained by rotating $\left(\bm{y}_{1},\dots,\bm{y}_{n}\right)$ 90 degrees clockwise about the $y_{i_{0}}-y_{j_{0}}$ plane with the center of $\Omega_{l}$ as the center of rotation, then for $\forall i=1,\dots,n$,
    $\widetilde{y}_{ii_{0}}=$ $\cos\left(-\frac{\pi}{2}\right)\left(y_{i,i_{0}}-\frac{a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}}{2}\right)-\sin\left(-\frac{\pi}{2}\right)\left(y_{i,j_{0}}-\frac{a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}}{2}\right)+\frac{a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}}{2}$
    $=$ $y_{i,j_{0}},$
    $\widetilde{y}_{ij_{0}}=$ $\sin\left(-\frac{\pi}{2}\right)\left(y_{i,i_{0}}-\frac{a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}}{2}\right)+\cos\left(-\frac{\pi}{2}\right)\left(y_{i,j_{0}}-\frac{a_{j_{0}}^{(l)}+b_{j_{0}}^{(l)}}{2}\right)+\frac{a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}}{2}$
    $=$ $a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}-y_{i,i_{0}},$
    and $\widetilde{y}_{ij}=y_{ij}$, for $j\neq i_{0}\,\text{and}\,j\neq j_{0}$. As is readily apparent,
    $\left|\mu_{j}^{(l)}-\widetilde{\hat{\mu}}_{j}^{(l)}\right|=\begin{cases}\left|\mu_{j}^{(l)}-\hat{\mu}_{j}^{(l)}\right|,&j\neq i_{0}\,\text{and}\,j\neq j_{0},\\ \left|\mu_{j_{0}}^{(l)}-\hat{\mu}_{j_{0}}^{(l)}\right|,&j=i_{0},\\ \left|\mu_{i_{0}}^{(l)}-\hat{\mu}_{i_{0}}^{(l)}\right|,&j=j_{0},\end{cases}$ (22)
    $\left|\Sigma_{jj}^{(l)}-\widetilde{\hat{\Sigma}}_{jj}^{(l)}\right|=\begin{cases}\left|\Sigma_{jj}^{(l)}-\hat{\Sigma}_{jj}^{(l)}\right|,&j\neq i_{0}\,\text{and}\,j\neq j_{0},\\ \left|\Sigma_{j_{0}j_{0}}^{(l)}-\hat{\Sigma}_{j_{0}j_{0}}^{(l)}\right|,&j=i_{0},\\ \left|\Sigma_{i_{0}i_{0}}^{(l)}-\hat{\Sigma}_{i_{0}i_{0}}^{(l)}\right|,&j=j_{0}.\end{cases}$ (23)
    When $i=i_{0}$ and $j=j_{0}$,
    $\left|\widetilde{\hat{\Sigma}}_{i_{0}j_{0}}^{(l)}\right|=$ $\left|\frac{1}{n}\sum_{k=1}^{n}\left(y_{k,j_{0}}-\frac{1}{n}\sum_{h=1}^{n}y_{h,j_{0}}\right)\left(a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}-y_{k,i_{0}}-\frac{1}{n}\sum_{h=1}^{n}(a_{i_{0}}^{(l)}+b_{i_{0}}^{(l)}-y_{h,i_{0}})\right)\right|$ (24)
    $=$ $\left|\frac{1}{n}\sum_{k=1}^{n}-\left(y_{k,j_{0}}-\frac{1}{n}\sum_{h=1}^{n}y_{h,j_{0}}\right)\left(y_{k,i_{0}}-\frac{1}{n}\sum_{h=1}^{n}y_{h,i_{0}}\right)\right|=\left|\hat{\Sigma}_{i_{0}j_{0}}^{(l)}\right|.$
    For other cases with $i\neq j$, it can also be similarly deduced that $\left|\widetilde{\hat{\Sigma}}_{ij}^{(l)}\right|=\left|\hat{\Sigma}_{ij}^{(l)}\right|$. Therefore, according to Eqs. (22)-(24), we have that the rotation transformation has no impact on the termination criterion (16). For Figure 3(a), the differences between the expectations and covariances of the two groups of particles and those of the uniform measure are
    $\left|\boldsymbol{\mu}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right|$ $=\left(0.25,0\right),\quad\left|\boldsymbol{\Sigma}^{(l)}-\boldsymbol{\hat{\Sigma}}^{(l)}\right|$ $=\begin{pmatrix}\frac{5}{48}&0\\ 0&\frac{1}{24}\end{pmatrix}$. (25)
    For Figure 3(b), the differences are
    $\left|\boldsymbol{\mu}^{(l)}-\hat{\boldsymbol{\mu}}^{(l)}\right|$ $=\left(0,0.25\right),\quad\left|\boldsymbol{\Sigma}^{(l)}-\boldsymbol{\hat{\Sigma}}^{(l)}\right|$ $=\begin{pmatrix}\frac{1}{24}&0\\ 0&\frac{5}{48}\end{pmatrix}$. (26)
    Obviously, Eqs. (25) and (26) merely involve a rearrangement of elements and thus does not affect the termination criterion (16). That is, the origin $\boldsymbol{0}$ in MSP is indistinctive, and there will be no such unreasonable phenomena of the star discrepancy in Figures 2 and 3.
4 Numerical experiments
This section conducts numerical experiments to test the performance of tree-based density estimation. Since a thorough comparative study between DSP and KDE has been performed in [18], we mainly focus on the performance of DSP-mix and MSP compared with DSP. In Eq. (5), the parameter $m$ specifying the number of candidate splitting points is set to $64$. To evaluate the performance of DSP, DSP-mix and MSP for high-dimensional datasets, we adopt two standard error metrics as in [18]. The first is the Kullback-Leibler (KL) divergence to quantify the similarity between two probability distributions,
$\textup{KL}(p^{\text{ref}}||p^{\text{num}})=\int p^{\text{ref}}(\bm{x})\log\left(\frac{p^{\text{ref}}(\bm{x})}{p^{\text{num}}(\bm{x})}\right)\mathrm{d}\bm{x},$ (27)
which can be evaluated by the arithmetic mean of the points $(\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{N})\sim p^{\text{ref}}$, namely,
$\textup{KL}(p^{\text{ref}}||p^{\text{num}})\approx\frac{1}{N}\sum_{i=1}^{N}\left(\log p^{\text{ref}}(\bm{x}_{i})-\log p^{\text{num}}(\bm{x}_{i})\right).$ (28)
The second metric is the Hellinger distance,
${H}(p^{\text{ref}},p^{\text{num}})=\frac{1}{2}\int\left(\sqrt{p^{\text{ref}}(\bm{x})}-\sqrt{p^{\text{num}}(\bm{x})}\right)^{2}\textup{d}\bm{x}=1-\int\sqrt{\frac{p^{\text{num}}(\bm{x})}{p^{\text{ref}}(\bm{x})}}p^{\text{ref}}(\bm{x})\textup{d}\bm{x},$ (29)
which can be approximated by
$\hat{H}(p^{\text{ref}},p^{\text{num}})\approx\hat{H}^{2}=1-\frac{1}{N}\sum_{i=1}^{N}\sqrt{\frac{p^{\text{num}}(\bm{x}_{i})}{p^{\text{ref}}(\bm{x}_{i})}}.$ (30)
For each setting of parameters, we run 10 simulations from independent sampling point sets and measure the standard derivation
$\textup{SD}=\sqrt{\frac{1}{10}\sum_{i=1}^{10}\left(\textup{KL}^{(i)}\right)^{2}-\left(\frac{1}{10}\sum_{i=1}^{10}\textup{KL}^{(i)}\right)^{2}},$ (31)
or
$\textup{SD}=\sqrt{\frac{1}{10}\sum_{i=1}^{10}\left(\hat{H}^{(i)}\right)^{2}-\left(\frac{1}{10}\sum_{i=1}^{10}\hat{H}^{(i)}\right)^{2}}.$ (32)
In practice, we record the maximal values of the standard derivations produced by DSP, DSP-mix and MSP (maxSD for short) to verify their robustness.
All simulations via C++ implementations run on the platform: AMD Ryzen 9950x (16 cores, 32 threads) with 256GB Memory (5600Mhz). The average computational time in different scenarios is recorded in Tables 2 and 4 (Beta mixtures), Tables 6 and 8(Gaussian mixtures) as well as Tables 10 and 12 (Cauchy mixtures). When the sample size is relatively small ($N\leq 10^{6}$), the computational costs for three methods are comparable. However, for large sample size ($N\geq 10^{7}$), both DSP-mix and MSP are evidently faster than DSP, and MSP even gains an increase in speed by a factor of two to twenty. Moreover, it is observed that the computational costs of both DSP-mix and MSP scale nearly linearly with respect to the sample size.
4.1 $d$-D Beta mixtures
First, we examine the $d$-D mixtures of betas,
$\bm{x}\sim\frac{1}{3}\left(\prod_{j=1}^{d}\operatorname{beta}(15,5)(x_{j})+\prod_{j=1}^{d}\operatorname{beta}(10,10)(x_{j})+\prod_{j=1}^{d}\operatorname{beta}(5,15)(x_{j})\right).$ (33)
Figure 5 visualizes this example for $d=2$. All three methods capture the main features ( three peaks) of the distribution.
(a) KL divergence ($d=2$) (left: DSP, middle: DSP-mix, right: MSP).
(b) Hellinger distance ($d=2$) (left: DSP, middle: DSP-mix, right: MSP).
Figure 4: 22-D Beta mixtures: The KL divergence and Hellinger distance under different NN and $\theta$.
In Figure 4, it is shown that the errors of the three methods are comparable, all decreasing as sample size $N$ increases. For fixed $N$, when the parameter $\theta$ decreases within a certain range, we can obtain finer partitioning and more accurate density estimation results. However, when $\theta$ goes too small, the error will increase instead, which is an indicator of the overfitting problem. Fortunately, the overfitting can be alleviated using a larger sample size.
(a) True density and samples.
(b) Adaptive partition (left: DSP, middle: DSP-mix, right: MSP).
(c) Density estimation (left: DSP, middle: DSP-mix, right: MSP).
Figure 5: 2-D Beta mixtures: Adaptive partitions and density estimators produced by DSP-mix, MSP and DSP with $N=1\times 10^{4}$ and $\theta=0.2$.
For high-dimensional tests with $d=15,20,30$, Figure 6 presents the convergence of error metrics with respect to sample size $N=10^{4},10^{5},10^{6},10^{7},10^{8}$, with the raw data collected in Tables 1 and 3. Figure 7 plots the partition level under different $d$, NN, with $\theta$ fixed to be $0.002$ with the raw data collected in Tables 2 and 4. From the above results, we make several observations.
  1. (1)
    All three methods are very robust as the standard derivations are much smaller than the mean errors.
  2. (2)
    The accuracy of all methods can be improved by increasing sample size $N$, which validates the theoretical error bound.
  3. (3)
    The accuracy of tree-based density estimation diminishes as dimension $d$ increases. This is because high-dimensional probability density tends to be concentrated in a very localized region (the typical set).
  4. (4)
    The performance of DSP and MSP is comparable, while MSP is significantly faster. DSP-mix works well for $d\leq 6$, but becomes less accurate as dimension $d$ increases. The reason is that the mixture discrepancy is a relaxation of the star discrepancy, so that the threshold $D^{\textup{mix}}\leq\frac{\theta\sqrt{N}}{n}$ is more readily to attain. As a result, the partition level decreases, which is confirmed in Figure 7.
  5. (5)
    The partition level $L$ increases as a smaller $\theta$ is chosen. When $N$ is large, it is seen that $L\sim\theta^{-1}$. When $\theta$ is too small, the partition ceases to be split further.
  6. (6)
    When $L$ is too large, the overpartitioning of space dramatically increases the computational cost but hampers the efficiency, due to the overfitting problem.
  7. (7)
    The overfitting problem can be alleviated by increasing sample size $N$.
(a) KL divergence (left: DSP, middle: DSP-mix, right: MSP).
(b) Hellinger distance (left: DSP, middle: DSP-mix, right: MSP).
(c) Partition level (left: DSP, middle: DSP-mix, right: MSP).
Figure 6: $d$-D Beta mixtures: Convergence of DSP, DSP-mix and MSP and the partition level $L$ with respect to NN for $d=15,20,30$.
(a) Partition level ($N=10^{4}$) (left: DSP, middle: DSP-mix, right: MSP).
(b) Partition level ($N=10^{5}$) (left: DSP, middle: DSP-mix, right: MSP).
(c) Partition level ($N=10^{6}$) (left: DSP, middle: DSP-mix, right: MSP).
(d) Partition level ($N=10^{7}$) (left: DSP, middle: DSP-mix, right: MSP).
Figure 7: $d$-D Beta mixtures: The partition level $L$ under different $\theta$, $N$ and $d$. It is seen that $L\sim\theta^{-1}$ when $N$ is sufficiently large.
Table 1: $d$-D Beta mixtures: The KL divergence and Hellinger distance under different $d$ and $\theta$. The sample size is fixed to be $N=10^{7}$.
$\theta$ $0.002$ $0.02$ $0.05$ $0.1$ $0.2$
$d$ Method KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$
2 DSP -0.0007 0.0011 0.0013 0.0008 0.0019 0.0008 0.0028 0.0010 0.0042 0.0013
DSP-mix 0.0001 0.0010 0.0017 0.0009 0.0029 0.0011 0.0047 0.0015 0.0087 0.0026
MSP 0.0001 0.0010 0.0011 0.0008 0.0016 0.0008 0.0023 0.0009 0.0033 0.0011
(maxSD) 0.0002 0.0001 0.0001 0.0000 0.0001 0.0001 0.0002 0.0001 0.0002 0.0001
4 DSP -0.2145 0.0542 0.0133 0.0075 0.0320 0.0105 0.0485 0.0140 0.0695 0.0192
DSP-mix -0.0546 0.0168 0.0246 0.0101 0.0501 0.0151 0.0802 0.0220 0.1201 0.0319
MSP -0.3167 0.0784 -0.0043 0.0076 0.0155 0.0078 0.0274 0.0096 0.0404 0.0123
(maxSD) 0.0012 0.0004 0.0002 0.0001 0.0006 0.0002 0.0008 0.0003 0.0012 0.0005
6 DSP -0.3601 0.0976 0.0914 0.0314 0.1490 0.0424 0.1978 0.0535 0.2531 0.0669
DSP-mix -0.0181 0.0328 0.1506 0.0441 0.2303 0.0614 0.3240 0.0822 0.4245 0.1053
MSP -0.4512 0.1166 0.0074 0.0291 0.0710 0.0306 0.1116 0.0362 0.1556 0.0445
(maxSD) 0.0003 0.0005 0.0010 0.0003 0.0015 0.0004 0.0028 0.0009 0.0052 0.0013
8 DSP -0.4151 0.1391 0.2459 0.0756 0.3636 0.0972 0.4529 0.1171 0.5552 0.1406
DSP-mix 0.0943 0.0677 0.3946 0.1043 0.5627 0.1387 0.7416 0.1738 0.8929 0.2052
MSP -0.4521 0.1458 0.0996 0.0670 0.2242 0.0752 0.1116 0.0362 0.3953 0.1043
(maxSD) 0.0008 0.0006 0.0029 0.0008 0.0058 0.0022 0.0130 0.0029 0.0126 0.0034
10 DSP -0.3763 0.1862 0.4781 0.1359 0.6679 0.1687 0.8025 0.1957 0.9590 0.2284
DSP-mix 0.2906 0.1210 0.7491 0.1838 1.0370 0.2353 1.2924 0.2779 1.4916 0.3158
MSP -0.3842 0.1875 0.2912 0.1211 0.4852 0.1386 0.6142 0.1582 0.7492 0.1838
(maxSD) 0.0015 0.0010 0.0048 0.0013 0.0126 0.0027 0.0175 0.0027 0.0165 0.0028
15 DSP 0.1017 0.3404 1.4157 0.3311 1.8129 0.3810 2.0849 0.4222 2.4313 0.4713
DSP-mix 1.1749 0.3092 2.0996 0.4180 2.7920 0.4909 3.1733 0.5313 3.5141 0.5784
MSP 0.1017 0.3406 1.1749 0.3097 1.5681 0.3443 1.8150 0.3765 2.0996 0.4179
(maxSD) 0.0099 0.0026 0.0220 0.0061 0.0326 0.0079 0.0286 0.0090 0.0345 0.0092
20 DSP 1.0627 0.5251 2.8794 0.5400 3.4996 0.5934 3.9360 0.6369 4.5112 0.6807
DSP-mix 2.6627 0.5242 4.1647 0.6455 5.2001 0.6993 5.7049 0.7391 6.2575 0.7851
MSP 1.0627 0.5257 2.6627 0.5237 3.2329 0.5605 3.6396 0.6003 4.1647 0.6452
(maxSD) 0.0195 0.0027 0.0213 0.0032 0.0205 0.0026 0.0269 0.0033 0.0295 0.0032
30 DSP 4.2883 0.8201 7.2776 0.8408 8.2751 0.8690 8.9656 0.8886 10.1624 0.9139
DSP-mix 7.1310 0.8337 9.8885 0.8990 11.3890 0.9258 12.1404 0.9391 12.9147 0.9523
MSP 4.2883 0.8195 7.1310 0.8342 8.0927 0.8590 8.7329 0.8761 9.8885 0.8989
(maxSD) 0.0382 0.0027 0.0289 0.0017 0.0417 0.0017 0.0296 0.0018 0.0252 0.0012
Table 2: $d$-D Beta mixtures: The partition level $L$ and computational time (in seconds) under different $d$ and $\theta$. The sample size is fixed to be $N=10^{7}$.
$\theta$ $0.002$ $0.02$ $0.05$ $0.1$ $0.2$
$d$ Method $L$ time $L$ time $L$ time $L$ time $L$ time
2 DSP 11601 19.34 4115 20.62 2969 27.54 2262 39.84 1634 66.47
DSP-mix 8874 15.35 4059 15.23 2915 15.75 2221 19.80 1590 31.96
MSP 8874 14.08 4342 14.31 3057 15.00 2334 14.74 1702 14.61
4 DSP 850435 947.49 28431 291.64 11696 292.27 6119 306.78 3275 329.77
DSP-mix 255326 41.38 25936 43.71 10381 62.42 5213 96.96 2619 165.91
MSP 1287088 49.38 65181 28.12 25332 27.64 14790 26.30 9079 25.90
6 DSP 1373682 1146.49 41757 430.30 17211 392.71 9272 393.10 5246 420.23
DSP-mix 270966 56.15 27402 62.56 11188 89.92 $56$ 140.38 2935 239.13
MSP 1731318 71.70 188819 42.25 76807 42.46 41359 40.14 22581 38.46
8 DSP 1707927 970.22 65024 595.31 26072 514.57 14138 489.84 8032 519.53
DSP-mix 285207 68.54 29847 80.34 12204 117.71 6260 179.82 3359 303.59
MSP 1845215 91.44 269574 59.93 108963 55.74 41359 40.14 29197 50.70
10 DSP 1921213 846.16 93104 751.74 36728 632.04 19901 585.76 11259 605.56
DSP-mix 303408 83.33 32637 96.91 13314 142.84 6924 216.97 3811 362.10
MSP 1949089 110.36 301774 72.19 122332 69.06 63136 65.12 32591 62.86
15 DSP 2235855 698.09 174225 1090.08 67961 875.88 36638 809.21 20170 859.77
DSP-mix 356114 121.08 40629 140.78 15801 204.62 8697 300.24 4900 507.12
MSP 2235891 166.61 356111 108.13 147379 100.36 77649 94.37 40629 91.47
20 DSP 2565816 723.45 268238 1436.26 105010 1174.55 56290 1083.46 30392 2503.68
DSP-mix 408780 160.17 47167 184.88 18688 260.86 10325 385.46 5456 652.46
MSP 2565816 226.91 408780 143.17 175550 134.04 91643 125.57 47167 121.13
30 DSP 3277027 898.88 425546 2079.65 169021 1626.90 95058 1825.46 45432 2000.92
DSP-mix 506018 236.90 57427 269.54 21036 379.71 11243 548.21 6627 914.78
MSP 3277027 359.31 506018 218.18 208819 196.87 116137 184.91 57427 178.54
Table 3: $d$-D Beta mixtures: The KL divergence and Hellinger distance under different $d$ and $N$. The parameter $\theta$ is fixed to be $0.002$.
NN $\times 10^{4} 1×1$$\times 10^{5}$ $\times 10^{6} 1×1$$\times 10^{7}$ $1\times 10^{8}$
$d$ Method KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$
15 DSP 2.0391 0.6561 1.2169 0.5513 1.4069 0.4060 0.1017 0.3404 0.4820 0.2105
DSP-mix 2.0391 0.6554 1.2169 0.5516 1.1811 0.4085 1.1749 0.3092 1.0136 0.2461
MSP 2.0391 0.6558 1.2169 0.5508 0.5660 0.4338 0.1017 0.3406 0.2728 0.2262
20 DSP 4.0528 0.8026 2.9441 0.7384 1.9095 0.6322 1.0627 0.5251 1.2661 0.3904
DSP-mix 4.0528 0.8027 2.9441 0.7382 2.7194 0.6178 2.6627 0.5242 2.2509 0.4465
MSP 4.0528 0.8030 2.9441 0.7382 1.9095 0.6320 1.0627 0.5257 1.0996 0.3998
30 DSP 9.0334 0.9385 7.6243 0.9253 5.8924 0.8828 4.2883 0.8201 3.6955 0.7627
DSP-mix 9.0334 0.9377 7.6243 0.9254 7.1100 0.8798 7.1310 0.8337 6.2093 0.7810
MSP 9.0334 0.9380 7.6243 0.9253 5.8924 0.8828 4.2883 0.8195 4.0776 0.7293
Table 4: $d$-D Beta mixtures: The partition level $L$ and computational time (in seconds) under different $d$ and $N$. The parameter $\theta$ is fixed to be $0.002$.
NN $\times 10^{4} 1×1$$\times 10^{5}$ $\times 10^{6} 1×1$$\times 10^{7}$ $1\times 10^{8}$
$d$ Method $L$ time $L$ time $L$ time $L$ time $L$ time
15 DSP 3289 0.14 28088 1.37 83373 161.28 2235855 698.09 5850719 29904.56
DSP-mix 3289 0.15 28088 1.47 122187 13.37 356114 121.08 1058493 1335.90
MSP 3289 0.15 28088 1.35 242887 15.29 2235891 166.61 10274742 1434.56
20 DSP 3782 0.20 32780 1.95 280051 20.97 2565816 723.45 8624426 37177.42
DSP-mix 3782 0.21 32780 2.09 143972 18.12 408780 160.17 1249503 1720.63
MSP 3782 0.21 32780 1.98 280051 21.12 2565816 226.91 11640399 1854.46
30 DSP $46$ 0.34 40468 3.27 352382 34.25 3277027 898.88 11265426 38853.64
DSP-mix $46$ 0.35 40468 3.61 183723 28.50 506018 236.90 1644069 2407.87
MSP $46$ 0.35 40468 3.44 352382 34.30 3277027 359.31 14813816 2909.42
4.2 $d$-D Gaussian mixtures
Now consider the $d$-D Gaussian mixtures [18],
$\bm{x}_{i}\sim\left(\sum_{i=1}^{4}\boldsymbol{\alpha}_{i}\mathcal{N}\left(\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}\right)\right),\,i=1,\dots,N,$ (34)
where $\boldsymbol{\alpha}=(0.4,0.3,0.2,0.1)$, $\boldsymbol{\mu}_{1}=0.3\mathbbm{1},\boldsymbol{\mu}_{2}=0.4\mathbbm{1},\boldsymbol{\mu}_{3}=0.5\mathbbm{1},\boldsymbol{\mu}_{4}=0.6\mathbbm{1}$, $\boldsymbol{\Sigma}_{1}=0.01\boldsymbol{I},\boldsymbol{\Sigma}_{2}=0.02\boldsymbol{I},\boldsymbol{\Sigma}_{3}=0.01\boldsymbol{I},\boldsymbol{\Sigma}_{4}=0.02\boldsymbol{I}$. Here $\mathbbm{1}$ is a $1\times d$ all-one vector, and $\boldsymbol{I}$ is a $d\times d$ identity matrix.
(a) KL divergence ($d=2$) (left: DSP, middle: DSP-mix, right: MSP).
(b) Hellinger distance ($d=2$) (left: DSP, middle: DSP-mix, right: MSP).
Figure 8: 22-D Gaussian mixtures: The KL divergence and Hellinger distance under different NN and $\theta$.
(a) True density and samples.
(b) Adaptive partition (left: DSP, middle: DSP-mix, right: MSP).
(c) Density estimation (left: DSP, middle: DSP-mix, right: MSP).
Figure 9: 2-D Gaussian mixtures: Adaptive partitions and density estimators produced by DSP-mix, MSP and DSP with $N=1\times 10^{4}$ and $\theta=0.2$.
Figure 9 presents the numerical solutions of DSP, DSP-mix and MSP. All of them possess the capability to capture the distribution structure. They feature fine grids at positions where the probability density changes significantly and sparse grids in unimportant regions.
From Figure 8, the V-shape curves with respect to $\theta$ are still observed in both error metrics. Therefore, it is suggested to avoid using too small parameter $\theta$ unless the sample size is large enough.
For high-dimensional tests under $d=15,20,30$, Figure 10 presents the convergence of error metrics with respect to sample size $N=10^{4},10^{5},10^{6},10^{7},10^{8}$, with the raw data collected in Tables 5 and 7. Figure 11 plots the partition level under different $d$, NN, with $\theta$ fixed to be $0.002$, with the raw data collected in Tables 6 and 8.
For the rapidly decreasing Gaussian distribution, the performance of DSP, DSP-mix and MSP is very similar to those for the localized Beta mixtures under different dimensions. It is still observed that the accuracy of all density estimators can be improved by increasing either sample size $N$ or increasing the partition level $L$. In terms of running time, DSP-mix and MSP are significantly faster and can generally achieve a tenfold or even twenty-fold speedup for large sample size ($N\geq 10^{7}$). Moreover, the trend $L\sim\theta^{-1}$ is still observed in all cases with only exception $d=2$.
(a) KL divergence (left: DSP, middle: DSP-mix, right: MSP).
(b) Hellinger distance (left: DSP, middle: DSP-mix, right: MSP).
(c) Partition level (left: DSP, middle: DSP-mix, right: MSP).
Figure 10: $d$-D Gaussian mixtures: Convergence of DSP, DSP-mix and MSP with respect to NN for $d=15,20,30$.
(a) Partition level ($N=10^{4}$) (left: DSP, middle: DSP-mix, right: MSP).
(b) Partition level ($N=10^{5}$) (left: DSP, middle: DSP-mix, right: MSP).
(c) Partition level ($N=10^{6}$) (left: DSP, middle: DSP-mix, right: MSP).
(d) Partition level ($N=10^{7}$) (left: DSP, middle: DSP-mix, right: MSP).
Figure 11: $d$-D Gaussian mixture: The partition level $L$ under different $\theta$, $N$ and $d$.
Table 5: $d$-D Gaussian mixtures: The KL divergence and Hellinger distance under different $d$ and $\theta$. The sample size is fixed to be $N=10^{7}$.
$\theta$ $0.002$ $0.02$ $0.05$ $0.1$ $0.2$
$d$ Method KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$
2 DSP -0.0008 0.0005 0.0014 0.0007 0.0020 0.0008 0.0030 0.0010 0.0042 0.0012
DSP-mix 0.0002 0.0009 0.0020 0.0009 0.0027 0.0010 0.0051 0.0016 0.0095 0.0028
MSP -0.0009 0.0005 0.0012 0.0007 0.0018 0.0007 0.0025 0.0009 0.0034 0.0011
(maxSD) 0.0001 0.0000 0.0001 0.0000 0.0001 0.0000 0.0002 0.0001 0.0003 0.0001
4 DSP 0.9405 0.0551 0.0046 0.0074 0.0187 0.0100 0.0356 0.0128 0.0510 0.0162
DSP-mix -0.0621 0.0173 0.0125 0.0092 0.0338 0.0128 0.0597 0.0195 0.0852 0.0250
MSP 0.8370 0.0798 -0.0118 0.0076 0.0062 0.0076 0.0164 0.0090 0.0262 0.0107
(maxSD) 3.4708 0.0026 0.0189 0.0026 0.0284 0.0026 0.0294 0.0026 0.0292 0.0027
6 DSP -0.3613 0.0969 0.0830 0.0284 0.1336 0.0368 0.1671 0.0437 0.2146 0.0547
DSP-mix -0.0253 0.0304 0.1213 0.0348 0.2003 0.0510 0.2791 0.0669 0.3264 0.0765
MSP -0.4532 0.1162 -0.0007 0.0271 0.0608 0.0273 0.0960 0.0308 0.1261 0.0355
(maxSD) 0.0003 0.0004 0.0008 0.0004 0.0023 0.0006 0.0029 0.0008 0.0034 0.0008
8 DSP -0.4183 0.1383 0.2277 0.0676 0.3186 0.0817 0.3812 0.0935 0.4916 0.1177
DSP-mix 0.0799 0.0626 0.3278 0.0832 0.5159 0.1182 0.6375 0.1377 0.7058 0.1510
MSP -0.4558 0.1455 0.0833 0.0623 0.2030 0.0667 0.2674 0.0734 0.3281 0.0831
(maxSD) 0.0011 0.0006 0.0025 0.0008 0.0036 0.0006 0.0047 0.0010 0.0058 0.0013
10 DSP -0.3841 0.1836 0.4510 0.1210 0.5888 0.1395 0.7002 0.1602 0.9045 0.1979
DSP-mix 0.2744 0.1115 0.6628 0.1523 0.9963 0.2029 1.1433 0.2224 1.2483 0.2418
MSP -0.3919 0.1848 0.2745 0.1113 0.4514 0.1217 0.5395 0.1313 0.6628 0.1523
(maxSD) 0.0021 0.0006 0.0036 0.0008 0.0061 0.0014 0.0122 0.0015 0.0118 0.0017
15 DSP 0.0593 0.3195 1.4166 0.2948 1.7380 0.3296 2.1080 0.3816 2.5918 0.4280
DSP-mix 1.2203 0.2818 2.2709 0.3894 2.8796 0.4337 3.1056 0.4576 3.5545 0.5211
MSP 0.0593 0.3198 1.2203 0.2817 1.5263 0.3001 1.7615 0.3296 2.2709 0.3894
(maxSD) 0.0053 0.0015 0.0081 0.0014 0.0205 0.0027 0.0264 0.0022 0.0295 0.0028
20 DSP 1.0081 0.4844 3.0613 0.4864 3.7281 0.5460 4.4446 0.5987 5.1962 0.6306
DSP-mix 2.8806 0.4746 4.9634 0.6092 5.6702 0.6371 6.1985 0.6827 7.3848 0.7647
MSP 1.0081 0.4843 2.8806 0.4748 3.3849 0.5061 4.0691 0.5638 4.9634 0.6094
(maxSD) 0.0200 0.0017 0.0311 0.0024 0.0483 0.0029 0.0591 0.0027 0.0668 0.0028
30 DSP 4.8871 0.7676 8.5323 0.7894 10.1693 0.8359 11.5031 0.8534 12.6868 0.8754
DSP-mix 8.3667 0.7800 12.4234 0.8571 13.8661 0.8963 15.4386 0.9284 17.0640 0.9359
MSP 4.8871 0.7679 8.3667 0.7805 9.9671 0.8262 11.3469 0.8446 12.4234 0.8570
(maxSD) 0.0320 0.0012 0.0418 0.0014 0.0426 0.0017 0.0555 0.0015 0.0612 0.0013
Table 6: $d$-D Gaussian mixtures: The partition level $L$ and computational time (in seconds) under different $d$ and $\theta$. The sample size is fixed to be $N=10^{7}$.
$\theta$ $0.002$ $0.02$ $0.05$ $0.1$ $0.2$
$d$ Method $L$ time $L$ time $L$ time $L$ time $L$ time
2 DSP 6603 7.88 3372 13.35 2302 18.66 1708 27.79 1239 45.49
DSP-mix 8353 6.92 3315 7.38 2260 9.03 1672 11.67 1200 20.86
MSP 7001 7.28 3608 8.83 2397 9.03 1788 8.99 1309 9.18
4 DSP 853559 882.76 27793 269.96 11255 270.44 5823 296.21 3155 323.50
DSP-mix 255266 21.33 25823 25.91 10390 47.30 5178 81.74 2649 149.73
MSP 1296208 32.90 62370 17.15 24249 14.46 14210 14.08 8777 13.24
6 DSP 1372339 1109.13 40997 426.41 16269 385.83 8894 398.37 4900 429.82
DSP-mix 271883 25.15 27674 36.71 11259 67.65 5568 118.32 2902 215.97
MSP 1732952 41.91 193079 26.18 78787 22.51 42202 20.70 23052 19.07
8 DSP 1711429 937.34 63630 613.12 25090 523.57 13821 512.66 7422 538.91
DSP-mix 287074 28.99 30227 47.96 11765 88.58 5933 152.11 3350 272.48
MSP 1850228 49.01 276904 34.20 110642 30.05 56644 27.22 29844 24.78
10 DSP 1927372 798.32 89458 774.72 35850 633.31 20018 606.81 10431 634.64
DSP-mix 301812 36.70 32765 58.30 12257 107.54 6529 181.35 3899 326.22
MSP 1955139 56.54 301284 38.56 119087 34.70 62357 31.48 32749 29.19
15 DSP 2268841 663.98 166721 988.71 72700 825.14 39392 851.34 19019 800.97
DSP-mix 336290 55.35 38372 85.67 14657 149.58 9097 243.25 5480 464.71
MSP 2268874 78.76 336290 54.82 143598 49.61 80693 45.85 38372 41.86
20 DSP 2721235 639.02 269423 1184.09 120788 1148.42 61092 1131.69 28893 914.53
DSP-mix 395153 77.47 42435 111.01 19350 185.18 12237 309.70 6016 613.64
MSP 2721235 105.96 395153 73.37 189409 66.72 96443 60.69 42435 55.44
30 DSP 3754332 734.10 526282 1923.58 212475 1829.85 99414 1466.28 50224 1479.35
DSP-mix 601142 115.55 59909 159.59 28074 265.94 13985 458.06 6604 843.57
MSP 3754333 167.07 601142 113.35 247671 98.94 114555 89.74 59909 85.51
Table 7: $d$-D Gaussian mixtures: The KL divergence and Hellinger distance under different dimensions $d$ and sample size $N$. The parameter $\theta$ is fixed to be $0.002$.
NN $\times 10^{4} 1×1$$\times 10^{5}$ $\times 10^{6} 1×1$$\times 10^{7}$ $1\times 10^{8}$
$d$ Method KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$
15 DSP 2.3652 0.6851 1.3628 0.5327 0.5726 0.4028 0.0593 0.3195 0.4881 0.2012
DSP-mix 2.3652 0.6853 1.3628 0.5332 1.1873 0.3716 1.2203 0.2818 0.9694 0.2203
MSP 2.3652 0.6848 1.3628 0.5331 0.5726 0.4033 0.0593 0.3198 0.2821 0.2201
20 DSP 4.9939 0.8507 3.5812 0.7394 2.0609 0.5949 1.0081 0.4844 1.2630 0.3555
DSP-mix 4.9939 0.8499 3.5812 0.7393 2.9840 0.5818 2.8806 0.4746 2.3519 0.4071
MSP 4.9939 0.8516 3.5812 0.7395 2.0609 0.5945 1.0081 0.4843 1.1115 0.3652
30 DSP 11.3623 0.9714 9.6601 0.9375 7.2038 0.8627 4.8871 0.7676 4.3272 0.6549
DSP-mix 11.3623 0.9717 9.6601 0.9374 9.1441 0.8652 8.3667 0.7800 7.3474 0.7180
MSP 11.3623 0.9713 9.6601 0.9376 7.2038 0.8626 4.8871 0.7679 4.2576 0.6557
Table 8: $d$-D Gaussian mixtures: The partition level $L$ and computational time (in seconds) under different $d$ and $N$. The parameter $\theta$ is fixed to be $0.002$.
NN $\times 10^{4} 1×1$$\times 10^{5}$ $\times 10^{6} 1×1$$\times 10^{7}$ $1\times 10^{8}$
$d$ Method $L$ time $L$ time $L$ time $L$ time $L$ time
15 DSP 3383 0.03 28693 0.37 246147 5.42 2268841 663.98 5838142 30802.07
DSP-mix 3383 0.03 28693 0.44 126396 5.11 336290 55.35 1042368 803.32
MSP 3383 0.03 28693 0.31 246147 5.49 2268874 78.76 10250539 752.33
20 DSP 3775 0.04 34852 0.57 305394 7.96 2721235 639.02 8503962 34682.44
DSP-mix 3775 0.04 34852 0.64 162165 6.76 395153 77.47 1170087 1076.32
MSP 3775 0.04 34852 0.48 305394 7.84 2721235 105.96 11362133 882.67
30 DSP 4433 0.07 46950 0.86 441887 13.93 3754332 734.10 14664912 42827.37
DSP-mix 4433 0.07 46950 1.04 217169 11.25 601142 115.55 1485038 1677.59
MSP 4433 0.07 46950 0.98 441887 13.49 3754333 167.07 15638938 1439.79
4.3 $d$-D Cauchy Mixtures
Finally, we consider a much more challenging problem: Learning the density of the multidimensional Cauchy mixtures [9]
$\bm{x}_{i}\sim\sum_{j=1}^{q}p_{j}C(\bm{x},\bm{m}_{j},\bm{u}_{j}),\quad C(\bm{x},\bm{m}_{j},\bm{u}_{j})=\prod_{k=1}^{d}\frac{u_{jk}}{\pi\left[u_{jk}^{2}+(x_{k}-m_{jk})^{2}\right]}.$ (35)
All the parameters are set to be: $q=4$, $p_{1}=p_{2}=p_{3}=p_{4}=0.25$, $u_{jk}=0.32$ for all $j,k$ and the centres $m_{1k}=-3$, $m_{2k}=-1$, $m_{3k}=1$, $m_{4k}=3$ for $k=1,\dots,d$.
(a) KL divergence ($d=2$) (left: DSP, middle: DSP-mix, right: MSP).
(b) Hellinger distance ($d=2$) (left: DSP, middle: DSP-mix, right: MSP).
Figure 12: 22-D Cauchy mixtures: The KL divergence and Hellinger distance under different NN and $\theta$.
(a) True density and samples.
(b) Heavy tail of Cauchy mixtures.
(c) Adaptive partition (left: DSP, middle: DSP-mix, right: MSP).
(d) Density estimation (left: DSP, middle: DSP-mix, right: MSP).
Figure 13: 2-D Cauchy mixtures: Adaptive partitions and density estimators produced by DSP-mix, MSP and DSP with $N=1\times 10^{4}$ and $\theta=0.2$.
One can sample the Cauchy distribution by using its cumulative function
$F\left(\frac{x}{u_{j}}\right)=\frac{1}{\pi}\left[\arctan\left(\frac{x}{u_{j}}-\frac{m_{j}}{u_{j}}\right)+\frac{\pi}{2}\right].$ (36)
As shown in Figure 13, the Cauchy distribution has a very heavy tail, although the typical set lies in a bounded domain $15,15]^{d}$. In fact, the Cauchy distribution has infinite mean and variance, so that direct Monte Carlo sampling may fail to converge. Therefore, we only try to reconstruct the density in the typical set $\Omega=[-15,15]^{d}$ and discard the particles outside $\Omega$.
Indeed, the heavy tail poses a significant challenge for the tree-based density estimation. Nevertheless, DSP, DSP-mix and MSP are able to capture the four-peak structure in the typical set, as clearly visualized in Figure 13. Figure 12 shows that the behavior of the KL divergence is a bit strange compared with Figures 4 and 8 for 2-D Beta mixtures and 2-D Gaussian mixtures. It may be also related to the heavy tail and we guess that, when the domain is enlarged, the performance of tree-based density estimator should be more sensitive to the partition level. As $\theta$ increases, the decision tree stops at a shallow level, producing coarser but more stable regions. This coarsening effectively suppresses the influence of isolated Cauchy outliers, causing the KL divergence to be smaller. That is, the KL divergence no longer serves as a satisfactory criterion for such heavy tail distribution as also pointed out in [32]. By contrast, the Hellinger distance always displays a similar behavior for the three mixtures.
For $d=15,20,30$, Figure 14 presents the convergence of error metrics with respect to sample size $N=10^{4},10^{5},10^{6},10^{7},10^{8}$. The raw data is collected in Tables 9 and 7. Figure 15 plots the partition level $L$ under different $d$, NN, with $\theta$ fixed to be $0.002$. The raw data is collected in Tables 10 and 12.
  1. (1)
    Overall, the errors diminish as either sample size $N$ and the partition level $L$ increase. The accuracy of MSP is comparable to DSP, whereas DSP-mix is less accurate as dimension $d$ increases.
  2. (2)
    Notably, for higher-dimensional cases with larger-sample, e.g., $d=30,N=10^{7}$, the KL divergence behaves smoothly across all $\theta$, confirming that the estimator can work well provided sufficient sample coverage of the heavy-tailed support.
  3. (3)
    The trend $L\sim\theta^{-1}$ is still observed in Figure 15. This gives us a simple way to estimate the partition level by first testing on a small dataset and using extrapolation, which also helps control the computational complexity.
(a) KL divergence (left: DSP, middle: DSP-mix, right: MSP).
(b) Hellinger distance (left: DSP, middle: DSP-mix, right: MSP).
(c) Partition level (left: DSP, middle: DSP-mix, right: MSP).
Figure 14: $d$-D Cauchy mixtures: Convergence of DSP, DSP-mix and MSP with respect to NN for $d=15,20,30$.
(a) Partition level ($N=10^{4}$) (left: DSP, middle: DSP-mix, right: MSP).
(b) Partition level ($N=10^{5}$) (left: DSP, middle: DSP-mix, right: MSP).
(c) Partition level ($N=10^{6}$) (left: DSP, middle: DSP-mix, right: MSP).
(d) Partition level ($N=10^{7}$) (left: DSP, middle: DSP-mix, right: MSP).
Figure 15: $d$-D Cauchy mixtures: The partition level $L$ under different $\theta$, $N$ and $d$.
Table 9: $d$-D Cauchy mixtures: The KL divergence and Hellinger distance under different $d$ and $\theta$. The sample size is fixed to be $N=10^{7}$.
$\theta$ $0.002$ $0.02$ $0.05$ $0.1$ $0.2$
$d$ Method KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$
2 DSP 0.4771 0.1892 0.3830 0.1789 0.3836 0.1789 0.3765 0.1795 0.3425 0.1806
DSP-mix 0.4229 0.1844 0.3832 0.1789 0.3841 0.1790 0.3777 0.1798 0.3454 0.1814
MSP 0.4414 0.1962 0.3825 0.1790 0.3832 0.1788 0.3755 0.1792 0.3407 0.1801
(maxSD) 0.1507 0.0002 0.0325 0.0001 0.0190 0.0001 0.0121 0.0001 0.0182 0.0001
4 DSP 0.1406 0.2198 0.1003 0.1922 0.1190 0.2019 0.0520 0.2128 0.1000 0.2255
DSP-mix 0.1276 0.1888 0.1334 0.2015 0.0730 0.2164 0.1408 0.2350 0.0086 0.2573
MSP 0.1231 0.2460 0.0724 0.1866 0.1129 0.1887 0.1199 0.1936 0.1312 0.2018
(maxSD) 0.2531 0.0001 0.0735 0.0001 0.0539 0.0001 0.0383 0.0001 0.0864 0.0003
6 DSP -0.1851 0.2715 -0.0588 0.2648 -0.1155 0.2925 -0.0748 0.3189 -0.0257 0.3481
DSP-mix -0.0882 0.2456 0.0001 0.3083 0.0902 0.3492 0.1615 0.3862 0.2793 0.4290
MSP -0.2477 0.2841 -0.1375 0.2455 -0.1319 0.2619 -0.0682 0.2813 -0.0062 0.3083
(maxSD) 0.1513 0.0001 0.1260 0.0003 0.1220 0.0004 0.1261 0.0005 0.0551 0.0007
8 DSP -0.6951 0.3909 -0.2798 0.3852 -0.1054 0.4298 0.0831 0.4677 0.1232 0.5088
DSP-mix -0.3676 0.3578 0.1077 0.4634 0.2991 0.5203 0.4819 0.5630 0.8625 0.6147
MSP -0.7145 0.3640 -0.3676 0.3578 -0.1418 0.3914 -0.0605 0.4254 0.1077 0.4634
(maxSD) 0.3285 0.0888 0.3285 0.0005 0.1064 0.0005 0.1113 0.0006 0.0973 0.0007
10 DSP -0.9253 0.4730 -0.0914 0.5246 0.1218 0.5767 0.4481 0.6182 0.5655 0.6593
DSP-mix -0.1747 0.4957 0.5718 0.6209 0.8718 0.6745 1.2397 0.7188 1.5495 0.7573
MSP -0.9281 0.4732 -0.1747 0.4958 0.0123 0.5399 0.2138 0.5793 0.5718 0.6209
(maxSD) 0.2714 0.0002 0.1065 0.0003 0.0482 0.0005 0.0852 0.0005 0.0401 0.0003
15 DSP -0.6121 0.7473 1.0220 0.8125 1.7181 0.8518 2.2469 0.8775 2.8725 0.9002
DSP-mix 0.9721 0.7955 2.6561 0.8834 3.4443 0.9143 4.0266 0.9337 4.5270 0.9455
MSP -0.6122 0.7473 0.9721 0.7955 1.6077 0.8333 2.1442 0.8597 2.6561 0.8834
(maxSD) 0.1532 0.0003 0.0943 0.0004 0.0908 0.0004 0.0768 0.0004 0.0234 0.0003
20 DSP 0.7989 0.9095 3.5479 0.9432 4.6367 0.9587 5.3676 0.9682 6.2838 0.9773
DSP-mix 3.5414 0.9382 6.1021 0.9721 7.1799 0.9818 7.9012 0.9856 8.4159 0.9892
MSP 0.7989 0.9095 3.5414 0.9382 4.5788 0.9533 5.2970 0.9629 6.1021 0.9721
(maxSD) 0.1970 0.0002 0.0908 0.0002 0.0681 0.0002 0.0884 0.0002 0.0258 0.0002
30 DSP 6.7636 0.9927 11.2454 0.9963 12.8170 0.9978 14.0692 0.9986 15.1920 0.9990
DSP-mix 11.2116 0.9960 15.1105 0.9988 16.4078 0.9992 17.1863 0.9996 18.0573 0.9998
MSP 6.7636 0.9927 11.2116 0.9960 12.7554 0.9975 13.9290 0.9984 15.1105 0.9988
(maxSD) 0.2761 0.0001 0.1189 0.0001 0.1874 0.0000 0.1519 0.0000 0.1898 0.0000
Table 10: $d$-D Cauchy mixtures: The partition level $L$ and computational time (in seconds) under different $d$ and $\theta$. The sample size is fixed to be $N=10^{7}$.
$\theta$ $0.002$ $0.02$ $0.05$ $0.1$ $0.2$
$d$ Method $L$ time $L$ time $L$ time $L$ time $L$ time
2 DSP 173841 151.72 16211 87.81 6639 105.19 3332 126.31 1691 146.15
DSP-mix 112188 8.84 16154 9.98 6583 16.92 3274 28.84 1625 52.87
MSP 280402 8.83 18032 6.07 6925 5.63 3698 5.25 2059 5.12
4 DSP 648039 687.78 27136 210.24 11910 192.20 6492 196.96 3659 205.85
DSP-mix 175907 14.01 17613 17.35 7307 29.68 3608 51.81 1890 94.56
MSP 1043302 17.84 115907 9.39 52032 9.07 28907 8.62 15741 8.29
6 DSP 1006568 678.35 53796 309.37 23027 257.94 12423 244.18 6906 245.08
DSP-mix 191134 18.10 19539 24.00 8038 42.36 4164 73.33 2158 137.66
MSP 1236975 22.90 188354 13.62 76132 12.40 38998 11.21 19520 10.63
8 DSP 1325701 553.24 85691 375.92 35788 290.85 18973 274.74 10181 272.09
DSP-mix 209324 23.47 21821 31.33 8940 54.64 4658 95.43 2379 177.23
MSP 1390910 29.26 209290 17.38 84843 16.10 43202 14.63 21821 13.77
10 DSP 1546736 461.04 119372 408.73 49117 319.62 25449 301.88 13155 296.88
DSP-mix 229815 28.18 23945 37.43 9809 66.84 5086 116.65 2651 210.02
MSP 1555407 33.94 229815 21.78 93768 18.77 47998 17.08 23945 16.16
15 DSP 1977162 340.83 204560 511.87 80332 421.72 39741 405.71 19590 395.28
DSP-mix 286280 38.17 28866 54.95 11644 97.07 5792 164.47 3199 290.12
MSP 1977232 51.67 286280 33.44 116620 28.26 58439 26.55 28866 24.54
20 DSP 2399337 328.32 279421 636.23 106726 576.62 51803 560.24 24388 534.52
DSP-mix 334965 53.23 32580 71.41 12502 126.16 6370 212.48 3633 363.47
MSP 2399363 71.01 334964 44.12 135152 37.84 67249 36.07 32580 32.42
30 DSP 3125241 386.49 389512 890.28 148772 890.74 68230 867.85 30332 839.55
DSP-mix 407694 81.72 35503 104.93 13466 182.92 7247 298.40 4264 513.47
MSP 3125248 109.90 407670 66.05 164124 59.87 77742 52.82 35503 48.01
Table 11: $d$-D Cauchy mixtures: The KL divergence and Hellinger distance under different $d$ and $N$. The parameter $\theta$ is fixed to be $0.002$.
NN $\times 10^{4} 1×1$$\times 10^{5}$ $\times 10^{6} 1×1$$\times 10^{7}$ $1\times 10^{8}$
$d$ Method KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$ KL $\hat{H}^{2}$
15 DSP 3.4022 0.9513 1.4427 0.9019 0.1989 0.8318 -0.6121 0.7473 1.0773 0.6183
DSP-mix 3.4022 0.9513 1.4427 0.9018 1.0771 0.8454 0.9721 0.7955 2.7906 0.7060
MSP 3.4022 0.9514 1.4427 0.9018 0.1739 0.8319 -0.6122 0.7473 2.7906 0.7062
20 DSP 7.1493 0.9904 4.3705 0.9767 2.4521 0.9500 0.7989 0.9095 2.7486 0.8506
DSP-mix 7.1493 0.9904 4.3705 0.9767 3.7805 0.9567 3.5414 0.9382 5.2577 0.9047
MSP 7.1493 0.9904 4.3705 0.9767 2.4521 0.9500 0.7989 0.9095 5.2577 0.9047
30 DSP 15.9541 0.9997 12.5253 0.9990 9.2726 0.9970 6.7636 0.9927 6.3239 0.9863
DSP-mix 15.9541 0.9997 12.5253 0.9990 11.4016 0.9977 11.2116 0.9960 10.3196 0.9940
MSP 15.9541 0.9997 12.5253 0.9990 9.2726 0.9970 6.7636 0.9927 10.3196 0.9940
Table 12: $d$-D Cauchy mixtures: The partition level $L$ and computational time (in seconds) under different $d$ and $N$. The parameter $\theta$ is fixed to be $0.002$.
NN $\times 10^{4} 1×1$$\times 10^{5}$ $\times 10^{6} 1×1$$\times 10^{7}$ $1\times 10^{8}$
$d$ Method $L$ time $L$ time $L$ time $L$ time $L$ time
15 DSP 2843 0.03 24527 0.25 222954 4.33 1977162 340.83 7840212 19084.37
DSP-mix 2843 0.03 24527 0.26 106704 3.53 286280 38.17 987999 596.05
MSP 2843 0.02 24527 0.24 222960 4.46 1977232 51.67 987999 570.70
20 DSP 3431 0.04 29100 0.37 272278 6.39 2399337 328.32 10550658 20532.97
DSP-mix 3431 0.04 29100 0.38 127445 4.97 334965 53.23 1086785 716.39
MSP 3431 0.04 29100 0.35 272278 6.36 2399363 71.01 1086785 732.24
30 DSP 4408 0.07 34327 0.67 351954 9.92 3125241 386.49 13079357 23887.23
DSP-mix 4408 0.07 34327 0.74 162494 7.88 407694 81.72 1129381 867.59
MSP 4408 0.07 34327 0.65 351954 10.32 3125248 109.90 1129381 894.90
5 Conclusion and discussion
This work proposes the DSP-mix and MSP methods, both of which adopt the same framework of the DSP method in [18] to learn a adaptive piecewise constant function to approximate the underlying probability density. The novelty lies in the measures of uniformity for observations. DSP-mix and MSP adopt the mixture discrepancy and moments, respectively, instead of the star discrepancy in DSP. The mixture discrepancy and moments are easier to compute and theoretically maintain reflection invariance and rotation invariance, which are the shortcomings of the star discrepancy. Numerical experiments up to 30 dimension demonstrate that MSP can maintain the same accuracy compared with DSP, while gaining an increase in speed by a factor of two to twenty for large sample size. DSP-mix achieves satisfactory accuracy and boosts the efficiency in low-dimensional tests ($d\leq 6$), but might lose accuracy in high-dimensional problems due to a reduction in partition level.
As a summary, the accuracy of DSP, DSP-mix and MSP can be improved by increasing the sample size $N$ under appropriate partition level $L$, although it still suffers from the overfitting problem when $L$ is too large. Fortunately, the overfitting can be alleviated as more samples are used. In practice, we can utilize the trend $L\sim\theta^{-1}$ to estimate the partition level by first testing on a small dataset and using extrapolation. This gives us a simple criterion to control the computational complexity.
There are still some issues worthy of further research, such as the quantitative error analysis of DSP-mix and MSP, the application to practical density estimation problems in higher dimensions, and more efficient uniformity measurement approaches.
Acknowledgement
This research was supported by the National Natural Science Foundation of China (Nos. 12325112, 12288101, 12571413) and the High-performance Computing Platform of Peking University. The authors are thankful to the anonymous referee for his/her valuable suggestions.